OpenStudy (sagewilson):

A is known to be 6,500 feet above sea level; AB = 600 feet. The angle at A looking up at P is 20°. The angle at B looking up at P is 35°. How far above sea level is the peak P? Find the height of the mountain peak to the nearest foot. Height above sea level __________________ = a0ft. (Hint: Draw a perpendicular from B to . Label the various angles. Compute BJ, then BP, then PQ, rounding to the nearest tenth in each step; and finally find 6,500 + PQ.)

OpenStudy (anonymous):

 may be the diagram is like this

OpenStudy (sagewilson):

maybe but i'm not 100% sure

OpenStudy (anonymous):

umm...it may take me a long time to do the steps... here comes my answer so far: tan35=PJ/BJ----(1) tan20=PQ/BJ----(2) BJ=PQ/tan20 tan35=PJ/(PQ/tan20) (tan35PQ)/tan20=PJ because PJ=PQ+600 therefore, (tan35PQ)/tan20=PQ+600 and solve PQ. after solving PQ, add 6500 and that's the height.

OpenStudy (anonymous):

it look a little bit messy :O if you don't understand, feel free to ask me